A beam of monochromatic light is incident from gla
Answer #1

n1sin(theta1) = n2sin(theta2)

where theta2 is angle of refraction.

case1) theta1 = 30 degrees

=> 1.5*sin(30 degrees) = 1*sin(theta2)

=> theta2 = 48.59 degrees

case2) theta1 = 60 degrees

=> 1.5*sin(60 degrees) = 1*sin(theta2)

=> sin(theta2) > 1

which is not possible => light will be reflected completely.

case3) theta1 = 25 degrees

=> 1.5*sin(25 degrees) = 1*sin(theta2)

=> theta2 = 39.27 degrees